# Formatting Instructions for ICLR 2025 Conference Submissions Residual Connections Harm Generative Representation Learning

### v3

@@ Abstract
We show that introducing a weighting factor to reduce the influence of identity shortcuts in residual networks significantly enhances semantic feature learning in generative representation learning frameworks, such as masked autoencoders (MAEs) and diffusion models. Our modification improves linear probing accuracy for both, notably increasing ImageNet accuracy from 67.8% to 72.7% for MAEs with a VIT-B/16 backbone, while also boosting generation quality for diffusion models. This significant gap suggests that, while residual connection structure serves an essential role in facilitating gradient propagation, it may have a harmful side effect of reducing capacity for abstract learning by virtue of injecting an echo of shallower representations into deeper layers. We ameliorate this downside via a fixed formula for monotonically decreasing the contribution of identity connections as layer depth increases. Our design promotes the gradual development of feature abstractions, without impacting network trainability. Analyzing the representations learned by our modified residual networks, we find correlation between low effective feature rank and downstream task performance.
@@ 1 Introduction
Residual networks (ResNets) (He et al., 2016) define a connection structure that has achieved near-universal adoption into modern architectures for deep learning. At the time of their development, supervised learning (e.g., ImageNet (Deng et al., 2009) classification) was the driving force behind the evolution of convolutional neural network (CNN) architectures. Residual networks solved a key issue: CNNs constructed of more than approximately 20 convolutional layers in sequence became difficult to train, leading to shallower networks outperforming deeper ones, unless additional techniques, such as auxiliary outputs (Szegedy et al., 2015) or batch normalization (Ioffe & Szegedy, 2015), were employed. Both ResNets, and their predecessor, highway networks (Srivastava et al., 2015a) provide elegant solutions to this trainability problem by endowing the network architecture with alternative shortcut pathways along which to propagate gradients. Highway networks present a more general formulation that modulates these shortcut connections with learned gating functions. However, given their sufficient empirical effectiveness, the simplicity of ResNet’s identity shortcuts (residual connections) makes them a preferred technique.
@@
While solving the gradient propagation issue, residual connections impose a specific functional form on the network; between residual connections, each layer (or block of layers) learns to produce an update slated to be added to its own input. This incremental functional form may influence the computational procedures learned by the network (Greff et al., 2017). Alternatives to residual and highway networks exist that do not share this functional form, but implement other kinds of skip-connection scaffolding in order to assist gradient propagation (Larsson et al., 2017; Huang et al., 2017; Zhu et al., 2018). Thus, shortcut pathways, rather than a specific form of skip connection, are the essential ingredient to enable the training of very deep networks. Nevertheless, nearly all modern large-scale models, including those based on the transformer architecture (Vaswani et al., 2017) incorporate the standard identity shortcut residual connection.
@@
This design choice holds, even as deep learning has shifted into an era driven by self-supervised training. The shift to self-supervision brings to the forefront new learning paradigms, including those based on contrastive (Wu et al., 2018; He et al., 2020; Chen et al., 2020; Caron et al., 2021; Grill et al., 2020), generative (Goodfellow et al., 2014; Karras et al., 2021; Ho et al., 2020; Song et al., 2021a; 2023; Rombach et al., 2022), and autoencoding (Kingma & Welling, 2013; He et al., 2022; Li et al., 2023) objectives. Many systems in the generative and autoencoding paradigms rely on “encoder-decoder” architectures, often styled after the original U-Net (Ronneberger et al., 2015), which contains additional long-range shortcuts between corresponding layers in mirrored symmetry about a central bottleneck. With representation learning as a goal, one typically desires that the middle bottleneck layer produce a feature embedding reflecting abstract semantic properties. The interaction of skip-connection scaffolding for gradient propagation with encoder-decoder architectures, self-supervised training objectives, and bottleneck representations has not been carefully reconsidered. This is a worrisome oversight, especially since, even in the supervised setting with standard classification architectures, prior work suggests that unweighted identity shortcuts may be a suboptimal design decision (Savarese & Figueiredo, 2017; Fischer et al., 2023).
@@
Intuitively, identity shortcut connections may not be entirely appropriate for capturing high-level, semantic representations as they directly inject low-level, high-frequency details of inputs into outputs, potentially compromising feature abstraction. We explore this issue within generative learning frameworks, including masked autoencoders (MAEs) (He et al., 2022) and diffusion models (Ho et al., 2020), leading paradigms for self-supervised image representation learning and generation. Our experiments demonstrate that identity shortcuts significantly harm semantic feature learning in comparison to an alternative we propose: gradually decay the weight of the identity shortcut over the depth of the network, thereby reducing information flow through it (Figure 2). With increasing layer depth, our approach facilitates a smooth transition from a residual to a feed-forward architecture, while maintaining sufficient connectivity to train the network effectively. Unlike prior work on learned gating (Srivastava et al., 2015a) or reweighting (Savarese & Figueiredo, 2017) mechanisms for residual connections, our method is a forced decay scheme governed by a single hyperparameter.
@@
A parallel motivation for our design stems from Huh et al. (2021), who show that features from residual blocks have higher rank than those produced by comparative feed-forward blocks. The smooth transition between residual and feed-forward behavior induced by our decay scheme regularizes deeper features toward exhibiting low-rank characteristics. Section 6 experimentally explores the correlation between our decayed identity shortcuts and low-rank feature representations. Figure 1 previews the corresponding improvements to representation learning. Our contributions are:
@@
We introduce decayed identity shortcuts, a simple architectural mechanism which enhances semantic feature abstraction in masked autoencoders and diffusion models.
@@
We identify a key correlation between our decayed identity shortcuts and low-rank inductive bias, empirically validating that our method improves classification accuracy and yields low-rank features with distinct cluster structures.
@@
Our novel design within an MAE yields a substantial performance boost in linear probing on ImageNet-1K (Deng et al., 2009) (72.7% from a baseline 67.8%).
@@
Ablation studies on ImageNet-100 show that smaller models equipped with decayed identity shortcuts outperform larger ones using standard residual connections. A VIT-S/16 model (Dosovitskiy et al., 2021) with our shortcut design outperforms a baseline VIT-B/16 (78.5% vs. 76.5%).
@@ 2 Related Work
At present, two families of approaches to self-supervised visual representation learning appear particularly promising: contrastive learning (Wu et al., 2018; He et al., 2020; Chen et al., 2020; Caron et al., 2021; Grill et al., 2020), which trains a discriminative model to maximize mutual information across image augmentations, and generative learning, via masked image modeling (Bao et al., 2022; He et al., 2022; Chen et al., 2024), which trains to reconstruct occluded pixels, or via diffusion denoising (Song et al., 2021b; Ho et al., 2020; Song et al., 2021a), which trains to reverse a process that mixes images with Gausssian noise. Some hybrid approaches (Zhou et al., 2022; Huang et al., 2023; Li et al., 2023) combine both families. Despite advancements, neither has demonstrated the same scalability (Singh et al., 2023) as seen in LLMs. This challenge is additional motivation for reconsidering the foundations of self-supervised network architectures.
@@
Low rank bias in neural networks. Over-parameterized neural networks exhibit surprising generalization capabilities, a finding seemingly in contradiction with classical machine learning theory (Neyshabur et al., 2019). This phenomenon implies the existence of some form of implicit regularization that prevents the model from overfitting. From the perspective of neural network parameterizations, Arora et al. (2019) suggest that linear models with more layers tend to converge to minimal norm solutions. In the context of CNNs, Huh et al. (2021) demonstrates that stacking more feed-forward layers compels the model to seek solutions of a lower rank, and Jing et al. (2020) reinforce this finding by adding more layers to an autoencoder’s bottleneck, thereby creating a representation bottleneck. In vision transformers, Geshkovski et al. (2024) examine the connection between attention blocks and mean-shift clustering (Cheng, 1995), showing that repeated attention operations result in low-rank outputs. Moreover, Dong et al. (2021) reveal that eliminating the shortcut connection from residual attention blocks causes features to degenerate to rank 1 structures doubly exponentially. From a different perspective, recent work (Radhakrishnan et al., 2022; Beaglehole et al., 2023; Radhakrishnan et al., 2024) shows training algorithms implicitly induce low-rank behavior in neural networks. Radhakrishnan et al. (2024) study the dimensionality reduction behavior of a recursive feature machine (Radhakrishnan et al., 2022) and effectively verify performance on low-rank matrix recovery.
@@ 3.1 Decayed Identity Shortcuts
where represents the output from the preceding layer, and denotes the transformation applied at the current layer. Although it is widely known that pure feed-forward architectures are susceptible to vanishing gradients when building deeper models, Huh et al. (2021) demonstrates that feed-forward modules offer implicit structural regularization, enabling deep models to generate abstract representations at bottlenecks.
@@
where is a rescaling factor to the residual path, controlling the information flow through the skip connection. Fully expanding this relation for a network with layers indexed from 0 to , we have that:
We see that the contribution of the input is scaled by each while each subsequent network block output omits scaling factors up to . Hence, the contribution of early features of the network is especially down-weighted, preventing the network from passing fine-grained detailed information to the bottleneck .
Decay schema. Rather than adopting a naive choice of as a constant across all layers, we choose to be a function parameterized by the layer index , where the contribution from the shortcut path is monotonically decreasing when increases:
@@ 3.2 Implementation Strategy
Skip connections for autoencoders. Since our method progressively decays the residual connections over network depth, it encourages the most abstract features to be learned by later layer. However, learning a highly abstract bottleneck is detrimental to the training objectives that aim for pixel-wise reconstruction, as they necessitate the preservation of information across all feature levels. To address this, we incorporate standard skip connections between the encoder and decoder, enabling the encoder to directly pass information from shallow layers to the decoder while learning increasingly abstract representations in the deeper encoder layers.
Stabilizing training with residual zero initialization. The model exhibits rapid feature norm growth at the beginning of training for . We suspect that the model learns to amplify the output feature norm of to counteract the significant decay applied to the residual connection. This growth leads to training instability and negatively impacts training convergence. To address this issue, we follow the implementation of previous works (Ho et al., 2020) and initialize the weights of the final output layer in each to zero instead of using the original Xavier uniform initialization (Glorot & Bengio, 2010). This approach significantly enhances training stability by limiting the rate of feature norm growth and enables us to explore training with even lower values of .
@@ 4 Experiments on Masked Autoencoder (MAE)
For masked autoencoders (MAEs) (He et al., 2022), we replace the residual connections in the encoder’s MLP and attention blocks with decayed identity shortcuts. The MAE operates by accepting images with a random subset of pixels masked out and learning to recover the discarded pixels. As Section 3.2 describes, we add skip connections between the encoder and decoder to facilitate learning abstract representations at the bottleneck. Since the original MAE has twice the number of encoder layers as decoder layers, we inject output from every other encoder layer into the corresponding decoder layer. To match spatial dimensions, injected encoder features are combined with learnable masked tokens before channel-wise concatenation. The implementation details for the training and evaluation are shown in Section A. He et al. (2022) show the desired representations appear at the end of encoder; we therefore apply our decaying schema only to the encoder.
@@ 4.1 Representation Learning on ImageNet-1k
We pre-train MAE on the ImageNet-1K train split (Deng et al., 2009) and follow recent works to evaluate the learned representations using linear probing and end-to-end finetuning, where we added a single linear head over learned representations to predict the image categories
For hyperparameters of pretraining MAE, including the learning rate schedule, total training epochs, and mask ratio, we follow the best settings found in the original paper. Please see the appendix for detailed experimental setups.
| Method | Objective | Augmentation | FT | LP |
| MoCo-v3(Chen et al., 2021) | Contrastive Loss | Full | 83.2 | 76.7 |
| DINO(Caron et al., 2021) | Contrastive Loss | Full | 83.3 | 78.2 |
| Con MIM(Yi et al., 2023) | Contrastive Loss | Affine + Mask | 83.7 | 39.3 |
| ADDP(VIT-L) (Tian et al., 2024) | Feature Loss (VQ) | Affine + Mask | 85.9 | 23.8 |
| Latent MIM(Wei et al., 2024) | Feature Loss | Affine+ Mask | 83.0 | 72.0 |
| Data2Vec(Baevski et al., 2022) | Feature Loss | Full+Mask | 84.2 | 68.0 |
| CAE(Chen et al., 2024) | Recon. + Feature Loss | Affine+Mask | 83.8 | 70.4 |
| I-JEPA(Assran et al., 2023) | Feature Loss | Affine+Mask | - | 72.9 |
| MAE(He et al., 2022) | Recon. Loss | Affine+Mask | 83.6 | 67.8 |
| Ours ( = 0.6) | Recon. Loss | Affine+Mask | 82.9 | 72.7 |
@@
We report the results in Table 1, where we show the linear probing performance of various self-supervised methods, which we categorize by their objectives, and data augmentation processes. In the top half of the table, we present methods that employ a contrastive loss. Although these methods produce the best probing accuracies, their success depends on a carefully designed data augmentation process, which may need to be tuned for each different data distribution. In the bottom half, we show several methods based on generative architecture, including ours, which do not rely on contrastive objectives. With the exception of Data2Vec, these methods only use a standard random affine data transformation (with random masking), which need not be distribution-reliant. Among these methods, MAE only uses a pixel-wise loss, I-JEPA , Latent MIM, and CAE use a latent feature alignment loss, and CAE uses both. Our method simply extends MAE by constructing an implicit feature bottleneck and shows significant improvements over the MAE baselines (72.7% vs. 67.3%), outperforming Data2Vec, Latent MIM and CAE and giving a probing accuracy competitive with I-JEPA, without needing to use explicit feature alignment.
@@
End-to-end fine-tuning, unlike linear probing which only trains a single linear layer, updates the entire network for image classification. Since the features can shift significantly from their pre-training state during end-to-end updating, we argue that this may not accurately reflect the quality of the learned representations. For example, DINO demonstrates superior performance in various downstream vision tasks compared to MAE, but its fine-tuning performance is worse than MAE. Similarly, ConMIM and ADDP exhibit poor linear probing performance, suggesting lower-quality representations, yet their fine-tuning performance surpasses that of contrastive learning methods. Nevertheless, we still provide the fine-tuning results for reference.
4.2 Embdding analysis.
We qualitatively evaluate the feature learning in Figure 3(a) by visualizing the t-SNE embeddings of the learned features for a random subset of test images. The embedding produced from our features displays much clearer separation than baseline, which has struggles to differentiate the categories.
To provide another qualitatively evlauation of our learned representations, we adopt the pixel-wise embedding approaches proposed by Zhang et al. (2024) to group the representations from the last layer of the encoder into a lower dimensional space. We use their default hyperparameters to cluster representations across images of COCO validation set. We visualize in Figure 4 the top 3 eigenvectors for both our approaches with and baseline. From the visualization, ours learns abstract representation and the object from the same categories have similar color, indicating a global consistent semantic grouping. The baseline MAE, on the other hand, doesn’t show clearly global semantic patterns overall and the low-level representations, e.g., edges, are also high-lightened, indicating a representations from multiple semantic hierarchies are entangled.
We further evaluate the clustering quantitatively, following the postprocessing protocol (Zhang et al., 2024) to produced unsupervised semantic segmentation and report the results as the mean intersection of union (mIoU). Ours (10.41 mIoU) achieve 6.31 mIoU improvement over baseline (4.10 mIoU), which support the qualitatively comparison.

### v4

# Residual Connections Harm Generative Representation Learning

@@ Abstract
We show that introducing a weighting factor to reduce the influence of identity shortcuts in residual networks significantly enhances semantic feature learning in generative representation learning frameworks, such as masked autoencoders (MAEs) and diffusion models. Our modification notably improves feature quality, raising ImageNet-1K K-Nearest Neighbor accuracy from 27.4% to 63.9% and linear probing accuracy from 67.8% to 72.7% for MAEs with a ViT-B/16 backbone, while also enhancing generation quality in diffusion models. This significant gap suggests that, while residual connection structure serves an essential role in facilitating gradient propagation, it may have a harmful side effect of reducing capacity for abstract learning by virtue of injecting an echo of shallower representations into deeper layers. We ameliorate this downside via a fixed formula for monotonically decreasing the contribution of identity connections as layer depth increases. Our design promotes the gradual development of feature abstractions, without impacting network trainability. Analyzing the representations learned by our modified residual networks, we find correlation between low effective feature rank and downstream task performance.
@@ 1 Introduction
Residual networks (ResNets) (He et al., 2016) define a connection structure that has achieved near-universal adoption into modern architectures for deep learning. At the time of their development, supervised learning (e.g., ImageNet (Deng et al., 2009) classification) was the driving force behind the evolution of convolutional neural network (CNN) architectures. Residual networks solved a key issue: CNNs constructed of more than approximately 20 convolutional layers in sequence became difficult to train, leading to shallower networks outperforming deeper ones, unless additional techniques, such as auxiliary outputs (Szegedy et al., 2015) or batch normalization (Ioffe & Szegedy, 2015), were employed. Both ResNets, and their predecessor, highway networks (Srivastava et al., 2015a) provide elegant solutions to this trainability problem by endowing the network architecture with alternative shortcut pathways along which to propagate gradients. Highway networks present a more general formulation that modulates these shortcut connections with learned gating functions. However, given their sufficient empirical effectiveness, the simplicity of ResNet’s identity shortcuts makes them a preferred technique.
@@
While solving the gradient propagation issue, residual connections impose a specific functional form on the network; between residual connections, each layer (or block of layers) learns to produce an update slated to be added to its own input. This incremental functional form may influence the computational procedures learned by the network (Greff et al., 2017). Alternatives to residual and highway networks exist that do not share this functional form, but implement other kinds of skip-connection scaffolding in order to assist gradient propagation (Larsson et al., 2017; Huang et al., 2017; Zhu et al., 2018). Thus, shortcut pathways, rather than a specific form of skip connection, are the essential ingredient to enable the training of very deep networks. Nevertheless, nearly all modern large-scale models, including those based on the transformer architecture (Vaswani et al., 2017) incorporate the standard residual connection.
@@
This design choice holds, even as deep learning has shifted into an era driven by self-supervised training. The shift to self-supervision brings to the forefront new learning paradigms, including those based on contrastive (Wu et al., 2018; He et al., 2020; Chen et al., 2020; Caron et al., 2021; Grill et al., 2020), generative (Goodfellow et al., 2014; Karras et al., 2021; Ho et al., 2020; Song et al., 2021a, 2023; Rombach et al., 2022), and autoencoding (Kingma & Welling, 2013; He et al., 2022; Li et al., 2023) objectives. Many systems in the generative and autoencoding paradigms rely on “encoder-decoder” architectures, often styled after the original U-Net (Ronneberger et al., 2015), which contains additional long-range shortcuts between corresponding layers in mirrored symmetry about a central bottleneck. With representation learning as a goal, one typically desires that the middle bottleneck layer produce a feature embedding reflecting abstract semantic properties. The interaction of skip-connection scaffolding for gradient propagation with encoder-decoder architectures, self-supervised training objectives, and bottleneck representations has not been carefully reconsidered. This is a worrisome oversight, especially since, even in the supervised setting with standard classification architectures, prior work suggests that unweighted identity shortcuts may be a suboptimal design decision (Savarese & Figueiredo, 2017; Fischer et al., 2023).
@@
Intuitively, identity shortcuts may not be entirely appropriate for capturing high-level, semantic features as they directly inject low-level, high-frequency details of inputs into outputs, potentially compromising feature abstraction. We explore this issue within generative learning frameworks, including masked autoencoders (MAEs) (He et al., 2022) and diffusion models (Ho et al., 2020), leading paradigms for self-supervised image representation learning and generation. Our experiments demonstrate that identity shortcuts significantly harm semantic feature learning in comparison to an alternative we propose: gradually decay the weight of the identity shortcut over the depth of the network, thereby reducing information flow through it (Figure 2). With increasing layer depth, our approach facilitates a smooth transition from a residual to a feed-forward architecture, while maintaining sufficient connectivity to train the network effectively. Unlike prior work on learned gating (Srivastava et al., 2015a) or reweighting (Savarese & Figueiredo, 2017) mechanisms for residual connections, our method is a forced decay scheme governed by a single hyperparameter.
@@
A parallel motivation for our design stems from Huh et al. (2021), who show that features from residual blocks have higher rank than those produced by comparative feed-forward blocks. The smooth transition between residual and feed-forward behavior induced by our decay scheme regularizes deeper features toward exhibiting low-rank characteristics. Section 6 experimentally explores the correlation between our decayed identity shortcuts and low-rank feature representations. Figure 1 previews the corresponding improvements to feature learning. Our contributions are:
@@
We introduce decayed identity shortcuts, a simple architectural mechanism which enhances feature abstraction in masked autoencoders and diffusion models.
@@
We identify a key correlation between our decayed identity shortcuts and low-rank inductive bias, empirically validating that our method improves classification accuracy and yields low-rank features.
@@
Our design within an MAE yields a substantial performance boost on ImageNet-1K (Deng et al., 2009): achieving a linear probing accuracy of 72.7% (up from a baseline of 67.8%) and a K-Nearest Neighbor accuracy of 63.9% (an improvement from the baseline of 27.4%).
@@
Smaller models with decayed identity shortcuts outperform larger ones using standard residual connections.
@@ 2 Related Work
At present, two families of approaches to self-supervised visual representation learning appear particularly promising. Contrastive representation learning (Wu et al., 2018; He et al., 2020; Chen et al., 2020; Caron et al., 2021; Grill et al., 2020) achieves state-of-the-art performance in most downstream classification tasks by training discriminative models to maximize mutual information between differently augmented views of images. However, these approaches rely on extensive and intricate data augmentation pipelines, necessitating domain expertise for adaptation to newe domains. Generative representation learning, via masked image modeling (Bao et al., 2022; He et al., 2022; Chen et al., 2024), which trains to reconstruct occluded pixels, or via diffusion denoising (Song et al., 2021b; Ho et al., 2020; Song et al., 2021a), which trains to reverse a process that mixes images with Gaussian noise, relying less on forming discriminative augmentations, learns to extract representations inherently along the generative process. Some hybrid approaches (Zhou et al., 2022; Huang et al., 2023; Li et al., 2023) combine both families. Despite advancements, neither has demonstrated the same scalability (Singh et al., 2023) as seen in LLMs. This challenge is additional motivation for reconsidering the foundations of self-supervised network architectures.
@@
Low rank bias in neural networks. Over-parameterized neural networks exhibit surprising generalization capabilities, a finding seemingly in contradiction with classical machine learning theory (Neyshabur et al., 2019). This phenomenon implies the existence of some form of implicit regularization that prevents the model from overfitting. From the perspective of neural network parameterizations, Arora et al. (2019) suggest that linear models with more layers tend to converge to minimal norm solutions. In the context of CNNs, Huh et al. (2021) demonstrate that stacking more feed-forward layers compels the model to seek lower rank solutions, and Jing et al. (2020) reinforce this finding by adding more layers to an autoencoder’s bottleneck, thereby creating a representation bottleneck. In vision transformers, Geshkovski et al. (2024) examine the connection between attention blocks and mean-shift clustering (Cheng, 1995), showing that repeated attention operations result in low-rank outputs. Moreover, Dong et al. (2021) reveal that eliminating the shortcut connection from residual attention blocks causes features to degenerate to rank 1 structures doubly exponentially. From a different perspective, recent work (Radhakrishnan et al., 2022; Beaglehole et al., 2023; Radhakrishnan et al., 2024) shows training algorithms implicitly induce low-rank behavior in neural networks. Radhakrishnan et al. (2024) study the dimensionality reduction behavior of a recursive feature machine (Radhakrishnan et al., 2022) and effectively verify performance on low-rank matrix recovery.
@@ 3.1 Decayed Identity Shortcuts
where represents the output from the preceding layer, and denotes the network block applied at the current layer. Although it is widely known that pure feed-forward architectures are susceptible to vanishing gradients when building deeper models, Huh et al. (2021) demonstrates that feed-forward modules offer implicit structural regularization, enabling deep models to generate abstract representations at bottlenecks.
@@
where is a rescaling factor to the residual path, controlling the information flow through the skip connection. Fully expanding this relation for a network with layers indexed from 1 to , we have that:
We see that the contribution of the input is scaled by each while each subsequent network block output omits scaling factors up to . Hence, the contribution of early features of the network is especially down-weighted, preventing the network from passing fine-grained detailed information to the bottleneck . During our experiments, we find the effective decay factor of the final layer, , plays a critical role in deciding the optimal decay rate when varying the network depth .
Decay schema. Instead of specifying as a constant across all layers, we choose to be a function parameterized by the layer index , where the contribution from the shortcut path is monotonically decreasing when increases:
@@ 3.2 Implementation Strategy
Skip connections for autoencoders. Since our method progressively decays the residual connections over network depth, it encourages the most abstract features to be learned by later layers. However, learning an abstract bottleneck is detrimental to the training objectives that aim for pixel-wise reconstruction, as they necessitate the preservation of detailed information. To address this, we incorporate standard skip connections between the encoder and decoder, enabling the encoder to directly pass information from shallow layers to the decoder while learning increasingly abstract representations in the deeper encoder layers.
Stabilizing training with residual zero initialization. The model exhibits rapid feature norm growth at the beginning of training for . We suspect that the model learns to amplify the output feature norm of to counteract the significant decay applied to the residual connection. This growth leads to training instability and negatively impacts training convergence. To address this issue, we follow the implementation of previous works (Ho et al., 2020) and initialize the weights of the final output layer in each to zero instead of using the original Xavier uniform initialization (Glorot & Bengio, 2010). This approach enhances training stability by controlling the growth of feature norm, especially with smaller .
@@ 4 Experiments on Masked Autoencoder (MAE)
For masked autoencoders (MAEs) (He et al., 2022), we replace the residual connections in the encoder’s MLP and attention blocks with decayed identity shortcuts. The MAE operates by accepting images with a random subset of pixels masked out and learning to recover the discarded pixels. Since the original MAE has twice the number of encoder layers as decoder layers, we build encoder-decoder skip connections by injecting output from every other encoder layer into the corresponding decoder layer. To match spatial dimensions, injected encoder features are combined with learnable masked tokens before channel-wise concatenation. The implementation details for the training and evaluation are shown in Section A. He et al. (2022) show the desired representations appear at the end of encoder; we therefore apply our decaying schema only to the encoder.
@@ 4.1 Representation Learning on ImageNet-1k
We follow the default hyperparameters from MAE (He et al., 2022) to pretrain ImageNet-1K train split (Deng et al., 2009) and use the standard protocol to evaluate the learned representation with end-to-end finetuning (FT), linear probing (LP) and K-Nearest Neighbour (KNN, K = 20), for image classification task. Please see the appendix for detailed experimental setups.
| Method | FT | LP | KNN |
| Contrastive representation learning | |||
| MoCo-v3(Chen et al., 2021a) | 83.2 | 76.7 | 66.6 |
| DINO(Caron et al., 2021) | 83.3 | 78.2 | 76.1 |
| Con MIM(Yi et al., 2023) | 83.7 | 39.3 | - |
| Generative representation learning | |||
| Data2Vec(Baevski et al., 2022) | 84.2 | 68.0 | 33.2 |
| I-JEPA(Assran et al., 2023) | - | 72.9 | - |
| CAE(Chen et al., 2024) | 83.8 | 70.4 | 51.4 |
| ADDP(VIT-L) (Tian et al., 2024) | 85.9 | 23.8 | - |
| Latent MIM(Wei et al., 2024) | 83.0 | 72.0 | 50.1 |
| MAE(He et al., 2022) | 83.6 | 67.8 | 27.4 |
| Ours ( = 0.6) | 82.9 | 72.7 | 63.9 |
| Feat. Dim. | Enc. Depth () | |||||||||
| 0.6 | 0.7 | 0.8 | 0.9 | 1.0 (Baseline) | (0, 1e-3) | [1e-3, 1e-2) | [1e-2, 1e-1) | [1e-1, 1]] | ||
| 384 | 12 | 78.5 | 78.1 | 75.2 | 73.5 | 69.2 | - | 78.5 | 78.1 | 75.2 |
| 768 | 12 | 83.6 | 81.8 | 79.8 | 79.2 | 76.5 | - | 83.6 | 81.8 | 79.8 |
| 1024 | 12 | 83.2 | 82.5 | 82.1 | 79.3 | 78.0 | - | 83.2 | 82.5 | 82.1 |
| 768 | 18 | 83.5 | 85.0 | 84.4 | 81.8 | 79.2 | 78.5 | 85.0 | 84.4 | 81.8 |
| 1024 | 24 | 84.3 | 86.0 | 84.5 | 84.3 | 81.4 | 82.4 | 86.0 | - | 84.3 |
| scheduler | UNet | LP | |
| Linear | Yes | 83.6 | |
| Linear | No | 61.5 | |
| Cosine | Yes | 82.8 | |
| Linear | Yes | 81.8 | |
| Cosine | Yes | 82.9 | |
| - | Learnable | Yes | 79.5 |
| Configurations | Decay Block | LP | |
| - | 76.5 | ||
| MLP & Atten. | 76.9 | ||
| MLP & Atten. | 82.6 | ||
| Atten. | 0.6 | 79.3 | |
| MLP | 0.6 | 80.6 | |
| MLP & Atten. | 0.6 | 83.6 | |
@@
We report the results in Table 1. In the top half of the table, we present methods that employ a contrastive loss. Although these methods produce the best probing accuracies, their success depends on a carefully designed data augmentation process, which may need to be tuned for each different data distribution. In the bottom half, we show several methods based on generative architecture. Our method simply extends MAE by constructing an implicit feature bottleneck and shows significant improvements over the MAE baselines for both linear probing (72.7% vs. 67.3%) and K-Nearest Neighbour (63.9% vs. 27.4%). outperforming Data2Vec, Latent MIM and CAE and giving a probing accuracy competitive with I-JEPA, without needing to use explicit feature alignment.
@@
End-to-end fine-tuning (FT), unlike linear probing which only trains a single linear layer, updates the entire network for image classification. Since the features can shift significantly from their pre-training state during end-to-end updating, we argue that this may not accurately reflect the quality of the learned representations. For example, DINO demonstrates superior performance in various downstream vision tasks compared to MAE, but its fine-tuning performance is worse than MAE. Similarly, ConMIM and ADDP exhibit poor linear probing performance, suggesting lower-quality representations, yet their fine-tuning performance surpasses that of contrastive learning methods. Nevertheless, we still provide the fine-tuning results for reference.